OpenStudy (anonymous):

Find the rectangular coordinates of the point with the polar coordinates ordered pair negative 7 comma 5 pi divided by 3.

OpenStudy (anonymous):

r^2=x^2+y^2 θ=tan^−1(y/x)

OpenStudy (misssunshinexxoxo):

:O -7 redoing it

OpenStudy (anonymous):

options: ordered pair 7 divided by 2 comma negative 7 divided by 2 ordered pair 7 square root 3 divided by 2 comma negative 7 divided by 2 ordered pair negative 7 divided by 2 comma 7 divided by 2 ordered pair negative 7 divided by 2 comma 7 square root 3 divided by 2

OpenStudy (misssunshinexxoxo):

Works out to the last option

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

I got a few more questions @misssunshinexxoxo

OpenStudy (misssunshinexxoxo):

Ok please post them as new questions and tag me

OpenStudy (anonymous):

alright